Lemon drops and jelly beans are mixed to make a 100 pound mix. If the lemon drops cost $1.90 per pound and jelly beans cost $1.20 per pound, how many pounds of jelly beans are needed to make the mix cost $1.48 per pound
Answer by stanbon(75887)

Lemon drops and jelly beans are mixed to make a 100 pound mix.
---
Quantity Eq::: L + J = 100 lbs
-----------
If the lemon drops cost $1.90 per pound and jelly beans cost $1.20 per pound, how many pounds of jelly beans are needed to make the mix cost $1.48 per pound
---
Value Eq::: 190L + 120J = 148*100 cents
-------------------------------------
Multipoly thru Quantity Eq by 190.
190L + 190J = 190*100
----
Subtract the Value Eq from that and solve for "J":
70J = 42*100
---
J = 60 lbs. (amt. of Jelly beans needed for the mix)
----
L + 60 = 100
L = 40 lbs. (amt. of Lemon drops needed for the mix)
=====================================================
Cheers,
Stan H.
=========